OpenStudy (anonymous):

Work-Out Corner has 5 more than 3 times as many exercise bicycles as The Gym. Together they have 21 bicycles. Solve the equation x + 3x + 5 = 21 to find the number of bicycles at Work-Out Corner.

Vocaloid (vocaloid):

start by combining like terms

Vocaloid (vocaloid):

x + 3x = ?

OpenStudy (anonymous):

1x + 3x is what? (its just 1+3 because the x makes it where 1x and 3x are equal.)

OpenStudy (anonymous):

We cannot just give you the answer. You need to work for it. Got it? If you try, you can do this.

OpenStudy (anonymous):

@southernbell19 Are you going to answer?

OpenStudy (anonymous):

I thought it was 4

Vocaloid (vocaloid):

almost hint: 1 + 3 = 4 x + 3x = __ fill in the blank

OpenStudy (anonymous):

4x

OpenStudy (anonymous):

Yes! so 4x + 5 = 21. Subtract 5. 4x + 5 - 5 = 21 - 5 what is 5-5? what is 21-5?

OpenStudy (anonymous):

5-5=0, 21-5=16

OpenStudy (seneca1):

The answer is $ but start by combining like terms so its x + 3x +4x but 21 - 5 4x = 16 4x/16 = 4